OpenStudy (anonymous):

The height of a pyramid is 3 meters shorter than the side of the base. The volume is 432m^3. Use V=1/3bh to fide the height. a 5 b18 c 9 d 12

OpenStudy (anonymous):

V = 1/3 * b * h b = (h+3)*(h+3) b = h^2+6h+9 V = 1/3*(h^2+6h+9)*h V = 1/3* (h^3+6h^2+9h) 432*3 = h^3+6h^2+9h 1296=h^3+6h^2+9h h = 9